. (a) Plot the following points in the same polar coordinates system



(3,π/4),(−3,π/4),(3,−π/4),(−3,−π/4)




(b) Convert into rectangular coordinates:



(4,−2π/3)





1
Expert's answer
2022-05-09T16:11:08-0400

(a)

"A(3,\u03c0\/4), B(\u22123,\u03c0\/4), C(3,\u2212\u03c0\/4), D(\u22123,\u2212\u03c0\/4)"




(b)


"(4\cos(-2\pi/3), 4\sin(-2\pi/3))"

"(-2, -2\sqrt{3})"